OpenStudy (anonymous):

Write an equation that would be solved by using multiplication. Then, show how to solve the equation.

OpenStudy (anonymous):

1/3x=7 3*1/3x=3*7 x=21

OpenStudy (anonymous):

Good job getusel:)

OpenStudy (anonymous):

dont forget to close question after it is helped with!

OpenStudy (anonymous):

....lol

OpenStudy (anonymous):

@hi8944 close the question if you don't need any further assistance!